OpenStudy (anonymous):

I fan and give medals I just need someone to check my answer:) Write a quadratic variation equation if g(x) varies directly with x2, and g(x) = 98 when x = 7. A.g(x) = 2x2 B.g(x) = ½x2 C.g(x) = 7x2 D.g(x) = 14x2 I think its C. but I may be wrong:)

Parth (parthkohli):

Umm, no. Again, use the old approach.

Parth (parthkohli):

In layman's language, all the options are pretty much same, but it is the number next to x^2 that differs.

OpenStudy (anonymous):

SO then its D.?

OpenStudy (anonymous):

because you x^2?

Parth (parthkohli):

And we can find it as well! Let's say that the number next to x^2 is c. Then we have the equation f(x) = cx^2. What does this become when x = 7? Just replace x with 7.

OpenStudy (anonymous):

49?

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Yeah, and the left side is given as 98. 98 = 49 * c Solve for c

OpenStudy (anonymous):

divide each side by 49?

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

so 98=c?

Parth (parthkohli):

No no... 98/49 = c

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so then the answer is a?

Parth (parthkohli):

98/49 = 2

Parth (parthkohli):

What do you think?

OpenStudy (anonymous):

yes thanks again for everything!